OpenStudy (anonymous):

Write (7x)^11/5 in radical form.

OpenStudy (anonymous):

\[(7x)^\frac{ 11 }{ 5 }\]

OpenStudy (mertsj):

Remember: top number is the power. Bottom number is the root.

OpenStudy (anonymous):

\[\sqrt[5]{ 7x^11 }\]?

OpenStudy (anonymous):

should be 7x^11, wouldn't type for some reason

OpenStudy (mertsj):

\[\sqrt[5]{7^{11}x ^{11}}\]

OpenStudy (mertsj):

\[\sqrt[5]{7^{10}(7)x ^{10}x}=7^2x^2\sqrt[5]7x{}=49x \sqrt[5]{7x}\]

OpenStudy (mertsj):

b

OpenStudy (anonymous):

i am the most confused kid.